OpenStudy (anonymous):

Find the value of 3 over the square root of 3.. A:3 B:3 times the square root of 3. C:the square root of 3. D:the square root of 3 over 3.

OpenStudy (anonymous):

\[3 \over \sqrt 3 \] = \[ 3 \over 3^\frac12\] Now, use the fact that : \[\frac{a^x}{a^y} = a^{x-y}\]

OpenStudy (matheducatormcg):

You can also rationalize the denominator. \[\frac{3}{\sqrt{3}}=\frac{3}{\sqrt{3}}\frac{\sqrt3}{\sqrt{3}}\] \[=\frac{3\sqrt{3}}{3}\] The 3's divide out and your result is....